The sum of two numbers is 98. The difference of the same two numbers is 40. What are the two numbers? Thanks!
Answer by mananth(16949)

let one number be x
the number be y
x+y = 98
x-y=40
add the two equations
2x=138
/2
x=69
first number 69
other number 98-69= 29